Page tree
Skip to end of metadata
Go to start of metadata

Attendees

Apologies

  • SWA

Agenda

      • Logs. Writing in temporary. STIX very big.
      • VM status - Engineering.
      • VM product status.
      • AOB

Discussion items

ItemNotes

Logs.

RC mentioned that

  1. The log of one of the SOC daemons is filling with observations that the SoloHI LLVM is repeatedly creating content in its temporary output directory (and then deleting it). This seems to be new(ish). If the VM needs to do this it is not a problem, and the SOC logging can be changed to ignore it.
  2. The STIX LLVM log is getting very very long.

NB After the meeting investigation by RC and LH showed that the reason for the first phenomenon above was repeated failed attempts by the SoloHI VM to process a specific request, but without any definitive failed output.

VM Status - engineering.

Emil indicated that he was planning to release a new VM version in the next month, just to incorporate small changes already made in the science pipeline.

For PHI, Martin said that a few minor changes were coming up, but that they were currently on the back burner. Additionally, work is ongoing on the decoding process. This will come in a new VM version, but is not imminent.

MK asked that RC look to see if old PHI archived logs were being deleted already. RC will respond directly.

LH mentioned an on-board problem which was expected to lead to problems for the LL pipeline, but which had already been fixed on board. She was keen to check if the same products were produced at the SOC as she was seeing from her own pipeline. One concern was the the VM might be reprocessing old inputs. These cross-checks were pursued after the meeting. RC asked about the smaller versions of the VMs. SoloHI confirmed that AL was continuing to work on this.

VM product status

Re STIX DW indicated that there are various issues to address, but that Shane has the information and things are converging.

Anik indicated that the products from SoloHI LLVM 00.01.24 can be ingested fine in the archive. Next was to see how the VM handled real data, and there should now be some after the checkout. So alongside or following the checks mentioned above of what output we are seeing in the SOC, Anik will follow up on the file content with Lynn.

AoB

We agreed that the next Telecon will be on the 30th of June.

Preparatory Actions

None.

New Actions

 - RC respond to MK re current state of PHI archived log deletion.